# RADLADS: Rapid Attention Distillation to Linear Attention Decoders at Scale

### v1

@@ 1 Introduction
acc. ratio (less random chance) Name Tokens Required Stages Lambada MMLU Others SUPRA (Mercat et al., 2024) 100B 1 0.913 0.216 0.869 LOLCatsb (Zhang et al., 2024a) 40M 2 - -0.288 0.635 MOHAWK (Bick et al., 2024b) 3-5B 3 0.938 -0.047 0.924 Mamba in the Llama (Wang et al., 2025) 20B 2 0.608 0.519 0.757 DiJiangb (Chen et al., 2024) 40B 1 - 0.720 - ARWKV (Yueyu et al., 2025) 60M/830M 2/3 0.893 0.801 0.935 Llamba (Bick et al., 2025) 8-12B 3 a0.951 a0.837 a0.987 RADLADS (ours) 350-700M 3 0.983 0.924 1.013
@@
While converting models, we found that pre-existing RWKV architectural choices were sometimes imperfectly matched to the needs of conversion. As a result, we developed two new architectures RAD-RWKV6 ("RADFinch") and RAD-RWKV7 ("RADGoose") based on RWKV-6 and RWKV-7 that allow conversion to proceed smoothly and the resultant model to be faster to inference than had we used the original RWKV-6 and RWKV-7 designs. We share these new architectural details, as well as some of the reasons behind the changes. We hope that other researchers may find this useful in converting transformers to their own linear attention designs.
@@
Two new simplified RWKV based architectures: RAD-RWKV6 ("RADFinch") and RAD-RWKV7 ("RADGoose"), used to convert softmax attention efficiently to linear attention.
Public release of several converted Qwen models: QRWKV6-7B ("Qwerky-7B-Base"), QRWKV6-7B-Instruct ("Qwerky-7B"), QRWKV7-7B-Instruct, QRWKV6-32B-Instruct ("Qwerky-32B"), QRWKV6-72B-Instruct ("Qwerky-72B"), and QRWKV6-QwQ-32B ("Qwerky-QwQ-32B").
@@ 2 Background
Other works such as Mamba in the Llama (Wang et al., 2025) have focused on a pipeline of progressive distillation followed by supervised fine-tuning (Kim & Rush, 2016) and direct preference optimization (Rafailov et al., 2024). Mamba in the Llama uses 20B tokens of training. However, it focuses on hybrid models rather than pure RNNs, and performs poorly when the softmax attention is removed completely.
A recent development in the conversion process, featured in LOLCats (Zhang et al., 2024a) and MOHAWK (Bick et al., 2024a) has been to separate the conversion process into two phases: attention alignment and full model knowledge distillation. The first phase aligns attention scores and/or attention hidden states with the teacher model, and the second performs traditional logit distillation. MOHAWK still requires training on a large amount of data. LOLCats requires only 40m tokens of training, but benchmark scores remain low, with MMLU accuracy lower than random guessing. To address this deficiency, they create a hybrid with full softmax attention in a sliding window (SWA), which improves scores but is no longer a purely recurrent model. We theorize that their low scores are due to the use of LoRAs for training, as well as vanilla linear attention rather than a more advanced recurrent architecture. This is most evident when examining their relative MMLU scores, which even with the inclusion of SWA remain low compared to ours.
@@ 4 Method
In RADLADS, our student model architecture is an exact copy of the teacher model architecture, with the exception of the attention blocks. These are replaced with a choice of purely recurrent attention-replacement time mixing blocks. In this paper we detail two such replacements, which are based on RWKV6-C2 and RWKV7, which we name RAD-RWKV6 ("RADFinch") and RAD-RWKV7("RADGoose") respectively. Our student models retain the runtime repetition of keys and values from Grouped Query Attention (GQA) (Ainslie et al., 2023) in the teacher model, if present. Depending on the attention-replacement chosen, we optionally include rotary positional embeddings (RoPE) (Su et al., 2023) from the teacher model. In cases where we include RoPE, it is applied at exactly the same point as in the teacher model.
@@ 4.4 Step 3: Context Length Extension
Name lmbda mmlu arc_c arc_e hella piqa winog Mistral-7B-v0.1-SUPRA 0.913 0.216 0.722 0.911 0.930 0.937 0.846 Mamba2-Llama3.0-8B-Instruct 0.608 0.519 0.724 0.868 0.901 0.903 0.391 MOHAWK-Phi1.5-1.3B 0.938 -0.047 0.830 0.968 0.936 0.959 0.927 DiJiang-Llama2.0-7B111Not Available: Neither the model weights nor the missing scores were published by the model authors. - 0.720 333The DiJiang paper incorrectly lists 0.403 and 0.561 for arc_c and arc_e respectively. We checked against Mercat et al. (2024)’s evals of this same teacher model and re-ran them here.0.973 333The DiJiang paper incorrectly lists 0.403 and 0.561 for arc_c and arc_e respectively. We checked against Mercat et al. (2024)’s evals of this same teacher model and re-ran them here.0.732 0.906 0.975 - Llamba-8B222Llamba-8B and QRWKV7-7B-Instruct-from72B are distilled from larger 70B and 72B models, respectively, but are compared with Llama3-8B and Qwen2.5-7B-Instruct in Table 3. 0.951 0.837 0.983 1.014 0.969 0.994 0.975 LOLCats-Llama3.0-8B111Not Available: Neither the model weights nor the missing scores were published by the model authors. - -0.029 0.534 0.864 0.750 0.886 0.143 ARWKV-7B 0.893 0.801 0.906 0.969 0.934 0.993 0.874 Our RADLADS models: QRWKV6-7B-Instruct 0.970 0.871 1.043 0.998 0.974 1.030 1.007 QRWKV6-7B-Instruct RoPE 0.969 0.880 1.054 1.003 0.971 1.018 0.981 QRWKV7-7B-Instruct 0.982 0.924 1.026 1.004 0.979 1.013 1.041 QRWKV7-7B-Instruct-from72B222Llamba-8B and QRWKV7-7B-Instruct-from72B are distilled from larger 70B and 72B models, respectively, but are compared with Llama3-8B and Qwen2.5-7B-Instruct in Table 3. 1.016 0.893 1.071 1.033 0.959 1.026 1.139 QRWKV6-32B-Instruct 0.986 0.909 1.066 1.037 0.964 1.005 1.235 QRWKV6-QwQ-32B 0.991 0.899 1.025 0.993 0.982 1.020 1.135 QRWKV6-72B-Instruct 1.004 0.899 1.015 1.008 0.974 0.968 1.123
@@
Name lmbda mmlu arc_c arc_e hella piqa winog Mistral-7B-v0.1-SUPRA 0.692 0.331 0.458 0.759 0.771 0.801 0.703 teacher: Mistral-7B-v0.1 0.758 0.624 0.538 0.809 0.810 0.821 0.740 Mamba2-Llama3.0-8B-Instruct 0.439 0.452 0.480 0.741 0.708 0.758 0.586 teacher: Llama3.0-8B-instruct 0.722 0.639 0.567 0.816 0.758 0.786 0.719 MOHAWK-Phi1.5-1.3B 0.501 0.242 0.441 0.740 0.602 0.755 0.717 teacher: Phi1.5-1.3B-base 0.534 0.418 0.480 0.756 0.626 0.766 0.734 DiJiang-Llama2.0-7B222Llamba-8B and QRWKV7-7B-Instruct-from72B are distilled from larger 70B and 72B models, respectively, but are compared with Llama3-8B and Qwen2.5-7B-Instruct in Table 3. - 0.407 0.427 0.626 0.694 0.775 - teacher: Llama2.0-7B - 0.468 333The DiJiang paper incorrectly lists 0.403 and 0.561 for arc_c and arc_e respectively. We checked against Mercat et al. (2024)’s evals of this same teacher model and re-ran them here.0.463 c0.764 0.740 0.782 - Llamba-8Ba 0.694 0.610 0.546 0.825 0.776 0.809 0.733 1st teacher: Llama3.1-8B-Instruct 0.730 0.680 0.551 0.817 0.793 0.811 0.739 2nd teacher: Llama3.1-70B-Instruct 0.757 0.823 0.624 0.867 0.847 0.837 0.787 LoLCATs-Llama3.0-8B222Llamba-8B and QRWKV7-7B-Instruct-from72B are distilled from larger 70B and 72B models, respectively, but are compared with Llama3-8B and Qwen2.5-7B-Instruct in Table 3. - 0.238 0.401 0.726 0.656 0.765 0.533 teacher: Llama3.0-8B 0.666 0.533 0.801 0.791 0.799 0.731 0.381 ARWKV-7B 0.622 0.624 0.522 0.797 0.768 0.792 0.687 QRWKV6-7B-Instruct 0.675 0.657 0.563 0.814 0.790 0.803 0.711 QRWKV6-7B-Instruct-RoPE 0.674 0.661 0.567 0.817 0.789 0.799 0.706 QRWKV7-7B-Instruct-RoPE 0.684 0.682 0.558 0.817 0.793 0.798 0.718 teacher: Qwen2.5-7B-Instruct 0.696 0.717 0.550 0.815 0.805 0.794 0.714 QRWKV7-7B-Instruct-from72B111Not Available: Neither the model weights nor the missing scores were published by the model authors. 0.707 0.667 0.572 0.833 0.782 0.801 0.739 teacher: Qwen2.5-72B-Instruct 0.751 0.834 0.632 0.859 0.874 0.836 0.763 QRWKV6-32B-Instruct 0.742 0.766 0.609 0.843 0.830 0.812 0.782 teacher: Qwen2.5-32B-Instruct 0.752 0.818 0.587 0.822 0.852 0.810 0.729 QRWKV6-QwQ-32B 0.662 0.743 0.564 0.806 0.830 0.804 0.732 teacher: Qwen2.5-QwQ-32B 0.668 0.799 0.556 0.810 0.841 0.798 0.705 QRWKV6-72B-Instruct 0.754 0.775 0.638 0.865 0.857 0.825 0.796 teacher: Qwen2.5-72B-Instruct 0.751 0.834 0.632 0.859 0.874 0.836 0.763
@@ 5 Language Modeling Performance
In Table 3 we compare pure RNN models output from various conversion methods across a set of standard benchmarks. These converted models were trained from variety of different teacher models and vary in terms of parameter count. Therefore, we compare them via ratio of benchmark accuracy scores between each student and teacher model, first subtracting the random chance accuracy for each benchmark. Negative ratios indicate that a model performed worse than random chance. Note that on nearly every benchmark, the RADLADS "Qwerky" models obtain higher score ratios than models converted via any other method. ("Qwerky" is our shorthand for "Qwen-RWKV-like") In Table 3 Qwerky6-72B demonstrates a new state-of-the-art in downstream performance for a pure RNN language model.
Next, in Tables 5 and 5 we show models converted to hybrid-attention variants, each containing some amount of softmax attention. Even when compared to these hybrid models that still employ softmax attention, the pure RNN Qwerky models show the highest MMLU ratios, and achieve similar or better ratios on other benchmarks.
@@ 6 Ablation Studies
We performed several ablation studies, adding or removing individual mechanisms to the RAD-RWKV6 architecture at the 7B scale. For the "use groupnorm" ablation we added the use of GroupNorm and removed the state balancing mechanism originally from RWKV6-C2. Results are shown in Table 6.
@@ 8 What Did Not Work
We originally included an initial "step 0", which was similar to step 1 but used a loss based on attention scores differences. For linear RNNs of the form
@@
This generalizes Mamba2, RWKV5, RWKV6, RWKV7, DeltaNet, and Gated DeltaNet, as well as others. See Table 1 in Peng et al. (2025) for a detailed list of such RNNs.
@@ B.1 RAD-RWKV6 Time Mixing
RAD-RWKV6 is a customized variation of RWKV6-C2 ("Finch-C2") (Goldstein et al., 2024), featuring a Gated Linear Attention (Yang et al., 2024) kernel (so no bonus), and a sigmoid gate. The use of the state balancing technique from RWKV6-C2 enables us to remove state normalization, improving downstream performance. Removal of bonus did not harm downstream performance. We do not typically use RoPE in this formulation.

### v3

@@ 1 Introduction
Relative scorea Name Tokens Required Stages Lambada MMLU Others SUPRA (Mercat et al., 2024) 100B 1 0.913 0.216 0.869 LOLCats/Hedgehogb (Zhang et al., 2024a) 40M 2 - -0.288 0.635 MOHAWK (Bick et al., 2024b) 3-5B 3 0.938 -0.047 0.924 Mamba in the Llama (Wang et al., 2024) 20B 2 0.608 0.519 0.757 DiJiangb (Chen et al., 2024) 40B 1 - 0.720 - ARWKV (Yueyu et al., 2025) 60M/830M 2/3 0.893 0.801 0.935 Llamba (Bick et al., 2025) 8-12B 3 c0.951 c0.837 c0.987 RADLADS (ours) 350-700M 3 0.983 0.924 1.013
@@
While converting models, we found that pre-existing RWKV architectural choices were sometimes imperfectly matched to the needs of conversion. As a result, we developed two new architectures RAD-RWKV6 (”RADFinch”) and RAD-RWKV7 (”RADGoose”) based on RWKV-6 and RWKV-7 that allow conversion to proceed smoothly and the resultant model to be faster to inference than had we used the original RWKV-6 and RWKV-7 designs. We share these new architectural details, as well as some of the reasons behind the changes. We hope that other researchers may find this useful in converting transformers to their own linear attention designs.
@@
Two new simplified RWKV based architectures: RAD-RWKV6 (”RADFinch”) and RAD-RWKV7 (”RADGoose”), used to convert softmax attention efficiently to linear attention.
Public release of several converted Qwen models: QRWKV6-7B-Base, QRWKV6-7B-Instruct, QRWKV7-7B-Instruct, QRWKV6-32B-Instruct, QRWKV6-72B-Instruct, and QRWKV6-QwQ-32B.
@@ 2 Background
Other works such as Mamba in the Llama (Wang et al., 2024) have focused on a pipeline of progressive distillation followed by supervised fine-tuning (Kim & Rush, 2016) and direct preference optimization (Rafailov et al., 2024). Mamba in the Llama uses 20B tokens of training. However, it focuses on hybrid models rather than pure RNNs, and performs poorly when the softmax attention is removed completely.
A recent development in the conversion process, featured in LOLCats (Zhang et al., 2024a) and MOHAWK (Bick et al., 2024a) has been to separate the conversion process into two phases: attention alignment and full model knowledge distillation. The first phase aligns attention scores and/or attention hidden states with the teacher model, and the second performs traditional logit distillation. MOHAWK still requires training on a large amount of data. LOLCats (Hedgehog version) requires only 40m tokens of training, but benchmark scores remain low, with MMLU accuracy lower than random guessing. To address this deficiency, they create a hybrid with full softmax attention in a sliding window (SWA), which improves scores but is no longer a purely recurrent model. We theorize that their low scores are due to the use of LoRAs for training, as well as vanilla linear attention rather than a more advanced recurrent architecture. This is most evident when examining their relative MMLU scores, which even with the inclusion of SWA remain low compared to ours.
@@ 4 Method
In RADLADS, our student model architecture is an exact copy of the teacher model architecture, with the exception of the attention blocks. These are replaced with a choice of purely recurrent attention-replacement time mixing blocks. In this paper we detail two such replacements, which are based on RWKV6-C2 and RWKV7, which we name RAD-RWKV6 (”RADFinch”) and RAD-RWKV7(”RADGoose”) respectively. Our student models retain the runtime repetition of keys and values from Grouped Query Attention (GQA) (Ainslie et al., 2023) in the teacher model, if present. Depending on the attention-replacement chosen, we optionally include rotary positional embeddings (RoPE) (Su et al., 2023) from the teacher model. In cases where we include RoPE, it is applied at exactly the same point as in the teacher model.
@@ 4.4 Step 3: Context Length Extension
Name lmbda mmlu arc_c arc_e hella piqa winog Mistral-7B-v0.1-SUPRA 0.913 0.216 0.722 0.911 0.930 0.937 0.846 Mamba2-Llama3.0-8B-Instruct 0.608 0.519 0.724 0.868 0.901 0.903 0.391 MOHAWK-Phi1.5-1.3B 0.938 -0.047 0.830 0.968 0.936 0.959 0.927 DiJiang-Llama2.0-7B111Not Available: Neither the model weights nor the missing scores were published by the model authors. - 0.720 333The DiJiang paper incorrectly lists 0.403 and 0.561 for arc_c and arc_e respectively. We checked against Mercat et al. (2024)’s evals of this same teacher model and re-ran them here.0.973 333The DiJiang paper incorrectly lists 0.403 and 0.561 for arc_c and arc_e respectively. We checked against Mercat et al. (2024)’s evals of this same teacher model and re-ran them here.0.732 0.906 0.975 - Llamba-8B222Llamba-8B and QRWKV7-7B-Instruct-from72B are distilled from larger 70B and 72B models, respectively, but are compared with Llama3-8B and Qwen2.5-7B-Instruct in Table 3. 0.951 0.837 0.983 1.014 0.969 0.994 0.975 LOLCatsHedgehog-Llama3.0-8B111Not Available: Neither the model weights nor the missing scores were published by the model authors. - -0.029 0.534 0.864 0.750 0.886 0.143 ARWKV-7B 0.893 0.801 0.906 0.969 0.934 0.993 0.874 Our RADLADS models: QRWKV6-7B-Instruct 0.970 0.871 1.043 0.998 0.974 1.030 1.007 QRWKV6-7B-Instruct RoPE 0.969 0.880 1.054 1.003 0.971 1.018 0.981 QRWKV7-7B-Instruct 0.982 0.924 1.026 1.004 0.979 1.013 1.041 QRWKV7-7B-Instruct-from72B222Llamba-8B and QRWKV7-7B-Instruct-from72B are distilled from larger 70B and 72B models, respectively, but are compared with Llama3-8B and Qwen2.5-7B-Instruct in Table 3. 1.016 0.893 1.071 1.033 0.959 1.026 1.139 QRWKV6-32B-Instruct 0.986 0.909 1.066 1.037 0.964 1.005 1.235 QRWKV6-QwQ-32B 0.991 0.899 1.025 0.993 0.982 1.020 1.135 QRWKV6-72B-Instruct 1.004 0.899 1.015 1.008 0.974 0.968 1.123
@@
Name lmbda mmlu arc_c arc_e hella piqa winog Mistral-7B-v0.1-SUPRA 0.692 0.331 0.458 0.759 0.771 0.801 0.703 teacher: Mistral-7B-v0.1 0.758 0.624 0.538 0.809 0.810 0.821 0.740 Mamba2-Llama3.0-8B-Instruct 0.439 0.452 0.480 0.741 0.708 0.758 0.586 teacher: Llama3.0-8B-instruct 0.722 0.639 0.567 0.816 0.758 0.786 0.719 MOHAWK-Phi1.5-1.3B 0.501 0.242 0.441 0.740 0.602 0.755 0.717 teacher: Phi1.5-1.3B-base 0.534 0.418 0.480 0.756 0.626 0.766 0.734 DiJiang-Llama2.0-7B222Llamba-8B and QRWKV7-7B-Instruct-from72B are distilled from larger 70B and 72B models, respectively, but are compared with Llama3-8B and Qwen2.5-7B-Instruct in Table 3. - 0.407 0.427 0.626 0.694 0.775 - teacher: Llama2.0-7B - 0.468 333The DiJiang paper incorrectly lists 0.403 and 0.561 for arc_c and arc_e respectively. We checked against Mercat et al. (2024)’s evals of this same teacher model and re-ran them here.0.463 c0.764 0.740 0.782 - Llamba-8Ba 0.694 0.610 0.546 0.825 0.776 0.809 0.733 1st teacher: Llama3.1-8B-Instruct 0.730 0.680 0.551 0.817 0.793 0.811 0.739 2nd teacher: Llama3.1-70B-Instruct 0.757 0.823 0.624 0.867 0.847 0.837 0.787 LoLCATsHedgehog-Llama3.0-8B222Llamba-8B and QRWKV7-7B-Instruct-from72B are distilled from larger 70B and 72B models, respectively, but are compared with Llama3-8B and Qwen2.5-7B-Instruct in Table 3. - 0.238 0.401 0.726 0.656 0.765 0.533 teacher: Llama3.0-8B 0.666 0.533 0.801 0.791 0.799 0.731 0.381 ARWKV-7B 0.622 0.624 0.522 0.797 0.768 0.792 0.687 QRWKV6-7B-Instruct 0.675 0.657 0.563 0.814 0.790 0.803 0.711 QRWKV6-7B-Instruct-RoPE 0.674 0.661 0.567 0.817 0.789 0.799 0.706 QRWKV7-7B-Instruct-RoPE 0.684 0.682 0.558 0.817 0.793 0.798 0.718 teacher: Qwen2.5-7B-Instruct 0.696 0.717 0.550 0.815 0.805 0.794 0.714 QRWKV7-7B-Instruct-from72B111Not Available: Neither the model weights nor the missing scores were published by the model authors. 0.707 0.667 0.572 0.833 0.782 0.801 0.739 teacher: Qwen2.5-72B-Instruct 0.751 0.834 0.632 0.859 0.874 0.836 0.763 QRWKV6-32B-Instruct 0.742 0.766 0.609 0.843 0.830 0.812 0.782 teacher: Qwen2.5-32B-Instruct 0.752 0.818 0.587 0.822 0.852 0.810 0.729 QRWKV6-QwQ-32B 0.662 0.743 0.564 0.806 0.830 0.804 0.732 teacher: Qwen2.5-QwQ-32B 0.668 0.799 0.556 0.810 0.841 0.798 0.705 QRWKV6-72B-Instruct 0.754 0.775 0.638 0.865 0.857 0.825 0.796 teacher: Qwen2.5-72B-Instruct 0.751 0.834 0.632 0.859 0.874 0.836 0.763
@@ 5 Language Modeling Performance
In Table 3 we compare pure RNN models output from various conversion methods across a set of standard benchmarks. These converted models were trained from variety of different teacher models and vary in terms of parameter count. Therefore, we compare them via relative score: the ratio of benchmark accuracy scores between each student and teacher model, first subtracting the random chance accuracy for each benchmark. Negative ratios indicate that a model performed worse than random chance. Note that on nearly every benchmark, the RADLADS ”QRWKV” models obtain higher score ratios than models converted via any other method. (”QRWKV” is our shorthand for ”Qwen with RWKV attention”) In Table 3 QRWKV6-72B-Instruct demonstrates a new state-of-the-art in downstream performance for a pure RNN language model.
Next, in Tables 5 and 5 we show models converted to hybrid-attention variants, each containing some amount of softmax attention. Even when compared to these hybrid models that still employ softmax attention, the pure RNN QRWKV models show the highest MMLU relative scores, and achieve similar or better relative scores on other benchmarks.
@@ 6 Ablation Studies
We performed several ablation studies, adding or removing individual mechanisms to the RAD-RWKV6 architecture at the 7B scale. For the ”use groupnorm” ablation we added the use of GroupNorm and removed the state balancing mechanism originally from RWKV6-C2. Results are shown in Table 6.
@@ 8 What Did Not Work
We originally included an initial ”step 0”, which was similar to step 1 but used a loss based on attention scores differences. For linear RNNs of the form
@@
This generalizes Mamba2 (Dao & Gu, 2024), RWKV5, RWKV6 (Peng et al., 2024), RWKV7 (Peng et al., 2025), DeltaNet (Schlag et al., 2021), and Gated DeltaNet (Yang et al., 2025), as well as others. See Table 1 in Peng et al. (2025) for a detailed list of such RNNs.
@@ 9 Conclusions
Acknowledgments
Eric Alcaide acknowledges support for this work from the Swiss State Secretariat for Education, Research and lnnovation (SERI) under contract number 23.00421.
@@ B.1 RAD-RWKV6 Time Mixing
RAD-RWKV6 is a customized variation of RWKV6-C2 (”Finch-C2”) (Goldstein et al., 2024), featuring a Gated Linear Attention (Yang et al., 2024) kernel (so no bonus), and a sigmoid gate. The use of the state balancing technique from RWKV6-C2 enables us to remove state normalization, improving downstream performance. Removal of bonus did not harm downstream performance. We do not typically use RoPE in this formulation.